A. An infinitely long wire carries a current of 3A determine the magnitude of the magnetic field 1.5 cm away.


B. What is the magnitude of the magnetic field 0.20 m away from a wire carrying a 5.00 A current? If the current has a vector direction out of the page (or screen), what is the direction of the magnetic field?


with solution thank you!




1
Expert's answer
2022-05-19T05:43:09-0400


A.

"B=\frac{\mu_0I}{2\pi r}"

"B=\frac{4\pi*10^{-7}*3}{2\pi *0.015}=4*10^{-5}\:\rm T"

B.

"B=\frac{4\pi*10^{-7}*5}{2\pi *0.20}=5*10^{-6}\:\rm T"

The direction of the magnetic field is perpendicular to the wire (counterclockwise).